Irving works for Sam's Food Market. He is paid weekly on the basis of a 40-hour work week at the rate of $9.50 an hour with time-and-a-half for overtime. During one week he worked 49 1/2 hours. What was the amount of his total earnings for the week?


1
Expert's answer
2022-04-07T11:51:02-0400

The amount of total earnings for the week:

"40\cdot 9.5+9.5\cdot9.5\cdot 1.5=\$\ 515.38"